Solve each system by the substitution method. Indicate wether each system is independent, inconsistent, or dependent.
x-y=3
3x-2y=3

Answer by tutorcecilia(2152)

x-y=3 [re-write]
-y=-x+3
y=x-3
.
substitute into the second equation:
3x-2y=3
3x-2(x-3)=3
3x-2x+6=3
x+6=3
x=3-6
x=-3
.
Substitute (x=-3) back into either of the original equatino and solve for the y-term:
3x-2y=3
3(-3)-2y=3
-9-2y=3
-2y=3+9
-2y=12
y=-6
.
Check by plugging (x=-3) and (y=-6) back into either equation and solve.